• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 00:02:42 on February 9,2009

    Entry number 259887

    keyword=shift summary - Feb 8 swing

    Start of shift: in controlled access, awaiting pass change.

    Problems with VDC left trip and moving spectrometer to 16 degrees.
    Jack comes in and resets VDC. Not clear why we could not, except
    that Jack is talented.

    One of bogie wheels spins near 44 degrees. Ed comes in and fixes it. Apparently they stuck paper under the wheel. It looks like this could be a continuing problem for the near future; Jack suggests we plan angle changes for during the day.

    ~8:10: access over.

    8:30: MCC calls and asks for BeO target.

    9:00: noticed funny light patterns on target. Eventually figure out is was tune beam + raster effect. Ask MCC to tune with raster off. JP says even BeO target is delicate.

    9:20: Ask MCC to tune beam through Compton.

    ~10:00: MCC asks where we want beam on target.

    10:30: ask MCC for 2 uA with nominal 2x2 raster.

    Been having regular MF2 alarms for much of last ~hour, which clear themselves within a few seconds. About 3 beeps typically.

    11:00 MCC asks for ABUs.

    11:17 MCC says awaiting authorization from radcom to raise ion chamber set points.

    Spend rest of shift determining that spot is not working on current
    run, then that beam is naturally a few mm wide without raster on,
    and needs to be better focused.